Status Report

SOHO Enters – and Emerges From – ESR (emergency sun reacquisition)

By SpaceRef Editor
January 17, 2001
Filed under ,

From SOHO’s Whats New page

January 16, 2001

    SOHO is in Normal Mode as of 20:49 UT. The last thruster activity was
    completed at about 19:50 UT. Some instrument checks and recovery
    activities were also performed today, but most of the instruments that
    were in safe mode because of the ESR will be returned to observation mode
    tomorrow or Thursday.

January 15, 2001

    The FOT and spacecraft engineers brought SOHO back into coarse roll
    pointing (CRP) mode at 17:49 UT on Monday, 2001
    January 15, and obtained star maps with the star sensor unit (SSU) that
    allowed them to determine the spacecraft roll orientation
    (-55.89 deg). With an accurate roll, a high gain antenna (HGA) pointing table
    was assembled, and medium-rate telemetry
    re-established through the HGA.

    After a brief period of instrument commanding, the FOT will:

    • Transition the spacecraft to reaction motor wheels (RMW) mode
    • Perform a momentum management maneuver
    • Transition to CRP
    • Roll back to nominal.

    The spacecraft should remain in CRP overnight.

January 14, 2001

    SOHO has been in ESR (emergency sun reacquisition) mode since
    Sunday, 2001 January 14 at approximately 21:02:02 UT. The proximate
    cause of the ESR was an attitude control unit (ACU) reset. Flight
    Operations Team personnel are in communication with the
    spacecraft, and are gathering information on the status of the ACU at
    the time of the reset before a decision can be made on when to return
    to Coarse Roll Pointing (CRP) and normal mode.

    The cause of the attitude control unit (ACU) reset remains unknown after
    inspecting the telemetry values prior to the reset, and there are no
    indications of energetic particle events causing the upset. The present
    roll rate is +0.2 deg/sec. A status meeting at 00:15 UT decided to leave
    the spacecraft in ESR mode for the night and start the recovery tomorrow
    morning. Instrument safing activities were completed at about 00:45 UT.

SpaceRef staff editor.